# MACE: Leveraging Audio for Evaluating Audio Captioning Systems

### v1

@@ III Multimodal Audio-Caption Evaluation
The MACE metric utilizes the Contrastive Language-Audio Pre-training (CLAP) model to overcome limitations of text-only evaluation methods by generating embeddings for both audio and text, enabling evaluation that accounts for acoustic content alongside textual descriptions. MACE is computed through three components: (1) CLAP audio and text embeddings assess the relevance of the generated caption to the audio content, (2) text embeddings evaluate acoustic similarity between the generated and reference captions, and (3) a fluency error penalty weights the similarity scores to ensure grammatical accuracy. Additionally, MACE can assess caption quality without a reference by using two of the three components.
@@
MACE captures semantic similarity, acoustic relevance, and linguistic quality within a single metric. By utilizing CLAP’s multimodal capabilities, it overcomes the core limitation of existing metrics that overlook the audio signal, offering a more comprehensive and accurate evaluation of audio caption quality.
@@ V-A Evaluating MACE
To evaluate the effectiveness of MACE, we conducted comprehensive benchmarking experiments on two widely-used datasets: Clotho-Eval and AudioCaps-Eval. The task consists of evaluating a pair of captions and selecting the better caption of the two. One of the captions in each pair is the preferred caption based on the scores assigned by human evaluators. TThe primary metric for comparison is pair accuracy, which measures a metric’s ability to correctly identify the human-preferred caption in each pair. To better understand and analyze the results, the tests are categorized into four groups: HC (two correct human captions), HI (one correct and one incorrect human caption), HM (one human and one machine-generated caption), and MM (two machine-generated captions). We compare MACE against a diverse set of metrics such as BLEU@1, BLEU@4, METEOR, ROUGE-L, CIDEr, SPICE, SPICE+, ACES, SPIDEr, X-ACE, FENSE and CLAIRA (with GPT-4o). all metrics on two categories (HC and HI) and is only slightly behind CLAIREA on the other two metrics (HM MACEM) and the overall accuracy (79.0% compared to 79.7%). The performance of CLAIREA dMACEs on the ability of GPT-4o to identify which of the two captions in a given pair is better - information which it may have already seen during training. Instead of GPT-4o, if one uses Gemini v1.5 (pro) [31], the accuracy falls down to 77.4%. On AudioCaps-Eval, MACE outperforms other metrics across all categories. The slight drop in performance on Clotho-Eval compared to AudioCaps-Eval, can be attributed to differences in audio content duration (Clotho-Eval contains longer audio clips) and content type (AudioCaps-Eval contains more speech). Notably, MACE demonstrates substantial improvements over the widely used FENSE metric on both datasets: AudioCaps-Eval (75.7% to 79.0%) and Clotho-Eval (85.3% to 88.1%).
@@ V-C MACE as objective metric
MACE can also be used as an objective metric [32] which does not require ground truth or reference caption for computation. To achieve this, we compute MACE but skip computing and adding $\text{S}_{\text{text-text}}$ score in Fig 1. This is equivalent to $\text{MACE}_{\text{AT}}$ score in Section V-B and Table II. Compared with FENSE which utilizes reference caption, $\text{MACE}_{\text{AT}}$ does not utilize reference caption is comparable or only slightly underperforms FENSE. Specifically, $\text{MACE}_{\text{AT}}$ outperforms FENSE by 1.45% on Clotho-Eval and underperforms FENSE by 3.16% on AudioCaps-Eval. This makes MACE ideal for large-scale audio captioning evaluations [33, 34, 35] where annotations are unavailable, costly, or restricted due to privacy.
@@ V-D Ablation analysis of fluency detection
Our findings indicate that the optimal threshold for error detection should be set significantly higher (0.97) than the default of 0.9, implying that the BERT model requires greater confidence to accurately identify fluency errors in audio captioning. Additionally, the penalty coefficient should be much lower (0.3) than the default of 0.9. This need for a higher detection threshold and reduced penalty suggests that the BERT model’s predictions does not directly align with human judgments of caption quality for the task of audio captioning. This can stem from the nature of AAC, where minor grammatical errors are less impactful on overall quality.

### v2

@@ III Multimodal Audio-Caption Evaluation
The MACE metric utilizes the Contrastive Language-Audio Pre-training (CLAP) model to overcome limitations of text-only evaluation methods by generating embeddings for both audio and text, enabling evaluation that accounts for acoustic content alongside textual descriptions. MACE is computed through three components: (1) CLAP audio and text embeddings assess the relevance of the generated caption to the audio content, (2) text embeddings evaluate acoustic similarity between the generated and reference captions, and (3) a fluency error penalty ensures grammatical accuracy. Additionally, MACE can assess caption quality without a reference by using two of the three components.
@@
MACE captures semantic similarity, acoustic relevance, and linguistic quality within a single metric. By utilizing CLAP’s multimodal capabilities, it overcomes the core limitation of existing metrics that overlook the audio signal, offering a more comprehensive and accurate evaluation of the caption quality.
@@ V-A Evaluating MACE
To evaluate the effectiveness of MACE, we conducted comprehensive benchmarking experiments on two widely-used datasets: Clotho-Eval and AudioCaps-Eval. The task consists of evaluating a pair of captions and selecting the better caption of the two. One of the captions in each pair is the preferred caption based on the scores assigned by human evaluators. TThe primary metric for comparison is pair accuracy, which measures a metric’s ability to correctly identify the human-preferred caption in each pair. To better understand and analyze the results, the tests are categorized into four groups: HC (two correct human captions), HI (one correct and one incorrect human caption), HM (one human and one machine-generated caption), and MM (two machine-generated captions). We compare MACE against a diverse set of metrics such as BLEU@1, BLEU@4, METEOR, ROUGE-L, CIDEr, SPICE, SPICE+, ACES, SPIDEr, X-ACE, FENSE and CLAIRA (with Gemini-v 1.5 and GPT-4o). all metrics on two categories (HC and HI) and is only slightly behind CLAIREA on the other two metrics (HM MACEM) and the overall accuracy (79.0% compared to 79.7%). The performance of CLAIREA dMACEs on the ability of GPT-4o to identify which of the two captions in a given pair is better - information which it may have already seen during training. Instead of GPT-4o, if one uses Gemini v1.5 (pro) [31], the accuracy falls down to 77.4%. On AudioCaps-Eval, MACE outperforms other metrics across all categories. The slight drop in performance on Clotho-Eval compared to AudioCaps-Eval, can be attributed to differences in audio content duration (Clotho-Eval contains longer audio clips) and content type (AudioCaps-Eval contains more speech). Notably, MACE demonstrates substantial improvements over the widely used FENSE metric on both the datasets: AudioCaps-Eval (75.7% to 79.0%) and Clotho-Eval (85.3% to 88.1%).
@@ V-C MACE as objective metric
MACE can also be used as an objective metric [32] which does not require ground truth or reference caption for computation. To achieve this, we compute MACE but skip computing and adding $\text{S}_{\text{text-text}}$ score in Fig 1. This is equivalent to $\text{MACE}_{\text{AT}}$ score in Section V-B and Table II. Compared with FENSE which utilizes reference caption, $\text{MACE}_{\text{AT}}$ does not utilize reference caption is comparable or only slightly underperforms FENSE. Specifically, $\text{MACE}_{\text{AT}}$ outperforms FENSE by 1.45% on Clotho-Eval and underperforms FENSE by 3.16% on AudioCaps-Eval. This makes MACE ideal for large-scale audio captioning evaluations [33, 34, 35] where annotations are either unavailable, costly, or restricted due to privacy concerns for users.
@@ V-D Ablation analysis of fluency detection
Our findings indicate that the optimal threshold for error detection should be set significantly higher (0.97) than the default of 0.9, implying that the BERT model requires greater confidence to accurately identify fluency errors in audio captioning. Additionally, the penalty coefficient should be much lower (0.3) than the default of 0.9. This need for a higher detection threshold and reduced penalty suggests that the BERT model’s predictions do not directly align with human judgments of caption quality for the task of automated audio captioning. This can stem from the nature of AAC, where minor grammatical errors are less impactful on the overall quality of the generated caption.